OpenStudy (anonymous):

Is (-2, -5) a solution of the system, 3x + 2y = 4 and 2x - 5y = -29? (Points : 1) yes no

OpenStudy (the_fizicx99):

Plug in -2 for x into both equations and -5 for y into both equations. If it gives you a true statement, then it is.

OpenStudy (anonymous):

ok it didnt so i guess its false

OpenStudy (the_fizicx99):

Yup, 3x + 2y = 4 3(-2) + 2(-5) = 4 -6 + (-10) = 4 -16 = 4 <-- not true However, if the y was 5, 3(-2) + 2(5) = 4 -6 + 10 = 4 <-- true.

OpenStudy (anonymous):

yeah but its false right

OpenStudy (the_fizicx99):

yup